OpenStudy (anonymous):

solve: log 4th root of x=square root log x

OpenStudy (anonymous):

\[\log \sqrt[4]{x }=\sqrt{\log x} \]

OpenStudy (anonymous):

x=1 , x = e^16

OpenStudy (anonymous):

1/4(logx) = (logx)^(1/2)

OpenStudy (anonymous):

substitute u = logx and solve the quadratic ... get it ?

OpenStudy (anonymous):

you should have u/4 = √u

OpenStudy (anonymous):

where did you get the u from

OpenStudy (anonymous):

is just a substitution to make the eq solvable ... you've never done substitutions?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it's a common technique

OpenStudy (anonymous):

teach me it

OpenStudy (anonymous):

how would you solve this (xy)^2 + 2xy +1 = 0 ?

OpenStudy (anonymous):

if you substitute xy = m (letter doesn't matter ) you'll get m^2 +2m +1 =0 ... now this is easier to solve ..right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

you just have to ... sense ... when is appropriate ... but it is quite common

OpenStudy (anonymous):

just like in your problem ... without the substitution would be very difficult to solve it

OpenStudy (anonymous):

so you substitue the log for u?(or whtever letter)?

OpenStudy (anonymous):

log x = u ...yes

OpenStudy (anonymous):

after using some properties of logs

OpenStudy (anonymous):

so u^1/2?

OpenStudy (anonymous):

u/4 = √u

OpenStudy (anonymous):

easier

OpenStudy (anonymous):

i got it thannks bro

OpenStudy (anonymous):

np